# FACE: Better Understanding Large Language Model Capabilities by Dissociating Language and Cognition

### v1

@@ Abstract
Large language models (LLMs) are primarily evaluated by overall performance on various text understanding and generation tasks. However, such a paradigm fails to comprehensively differentiate the fine-grained language and cognitive skills, rendering the lack of sufficient interpretation to LLMs’ capabilities. In this paper, we present FACE, a framework for Fine-grAined and Cognition-grounded LLMs’ Capability Evaluation. Specifically, we formulate LLMs’ evaluation in a multi-dimensional and explainable manner by dissociating the language-related capabilities and the cognition-related ones. Besides, through extracting the intermediate reasoning from LLMs, we further break down the process of applying a specific capability into three sub-steps: recalling relevant knowledge, utilizing knowledge, and solving problems. Finally, FACE evaluates each sub-step of each fine-grained capability, providing a two-faceted diagnosis for LLMs. Utilizing FACE, we identify a common shortfall in knowledge utilization among models and propose a straightforward, knowledge-enhanced method to mitigate this issue. Our results not only showcase promising performance enhancements but also highlight a direction for future LLM advancements.
FACE: Better Understanding Large Language Model Capabilities by Dissociating Language and Cognition
Xiaoqiang Wang1,2, Bang Liu1,2 and Lingfei Wu3 1DIRO & Institut Courtois, Université de Montréal 2Mila - Quebec AI Institute 3Anytime.AI {xiaoqiang.wang, bang.liu}@umontreal.ca, lwu@anytime-ai.com
@@ 1 Introduction
To address this challenge, most prior studies have assessed the performance of LLMs on different tasks based on independent benchmarks from various dimensions Liang et al. (2022); Srivastava et al. (2023); Gao et al. (2023), such as knowledge Hendrycks et al. (2020) and reasoning Zellers et al. (2019). Recently, motivated by building LLM-based AI assistants, some studies propose highly curated benchmarks with instance-level fine-grained annotations, such as difficulty and required skills Mialon et al. (2023); Ye et al. (2023), for holistic evaluation of LLMs.
@@
However, the existing studies overlook a crucial distinction between language and cognition Monti et al. (2012); Blank et al. (2014), rendering the insufficient understanding of a model’s true capabilities and effectiveness in tasks. For instance, in the context of generative question answering, a model adept at extracting information but struggling to form a coherent understanding may exhibit similar overall performance to another model with profound cognitive insights but difficulties in articulating accurate responses. We argue that a multi-dimensional and interpretable understanding of LLMs’ capabilities can not only accurately unveil their inherent limitations, but also help us to better identify why one model outperforms the other and how the different capabilities correlate. Additionally, such insights allow us to provide tailored guidance to improve training efficiency or facilitate more advanced model development.
@@
In this paper, we propose FACE, a fine-grained capability evaluation framework for LLMs. Specifically, FACE dissociates the language-related and cognition-related capabilities of LLMs and organizing them into four distinct axes: Linguistic Knowledge, Formal Knowledge, World Modeling, and Social Modeling. This categorization is grounded in neuroscience evidence manifesting that language processing and cognitive processes, like memory and reasoning, operate differently in the brain. Drawing from this insight, we adapt a range of existing benchmarks into a unified question-answering format. We then develop specific instructions for each capability, allowing FACE to evaluate LLMs through a method known as few-shot instruction-following.
@@
Furthermore, we break down the application of a specific capability into three sub-steps: knowledge recall, knowledge utilization, and problem-solving, by iteratively drawing out the model’s intermediate reasoning. After evaluating each sub-step, FACE can reveal the quality of knowledge encoded in the model, and effectiveness in applying relevant knowledge to solve practical problems, offering a more comprehensive evaluation than a single performance metric could.
@@
Our findings reveal a notable gap in capabilities between open-source and proprietary models, especially for cognition-related capabilities. Additionally, we found that many models have difficulties in applying knowledge effectively. To address this, we suggest a knowledge-enhance remedy by incorporating relevant knowledge text as additional input. Experimental results show that it can help the backbone model achieve about 90% performance compared to its instruction-tuned counterparts.
@@
| Conducting word-based formal reasoning through understanding shallow semantics. | Skill: | numeric, logic, and manipulation. | |
@@ 2 Methodology
In this section, we introduce FACE framework, designed for fine-grained and cognition-grounded LLMs’ capability evaluation. Specifically, we first define the taxonomy for LLMs’ capabilities based on the distinction between language and cognition, which is drawn upon insights from neuroscience Fedorenko and Varley (2016); Mahowald et al. (2023). Based on this, we transform a variety of existing benchmarks into the unified question-answering format, design capability-specific instruction, and frame FACE via few-shot instruction-following. Furthermore, we break down the evaluation process for each capability into a three-step reasoning approach. This involves identifying the knowledge pertinent to the input, examining how the model applies this knowledge in practical contexts, and assessing the effectiveness of its problem-solving. By evaluating each of these steps, FAC²E provides a comprehensive overview of the model’s performance, offering a more nuanced understanding of LLMs’ intrinsic capabilities.
@@ 2.1 Formulation of LLMs’ Capabilities
Human language processing has been investigated for a long time in cognitive science and neuroscience, which robustly attribute language and reasoning to different brain areas, namely “language network” and “multi-demand network” Duncan (2010); Scott et al. (2017). The former is sensitive to linguistic regularities and formal operations, and damage to the language network leads to linguistic deficits. The latter responds actively to a broader range of cognitively demanding processes, such as reasoning and memory. Motivated by this separated relationship, we define the LLMs’ capabilities as a 4-dimensional schema as shown in Table 1. We then proceed to explain each capability and underline the specific skills they encompass.
@@
Linguistic Knowledge. For LLMs to effectively generate language, they must first grasp text at a basic linguistic level, which includes understanding both grammaticality and semantics. Grammaticality encompasses the rules that govern language structure, spanning from the sounds (phonological) and words (lexical) to the arrangement of words in phrases and sentences. To capture this grammatical structure as comprehensively as possible, especially given the challenges conventional models face in benchmarks like BLiMP Warstadt et al. (2020), we focus on four key skills: agreement (anaphor and subject-verb relationships), licensing (negative polarity items and reflexive pronouns), managing long-distance dependencies (filler-gap constructions and cleft sentences), and navigating garden-path sentences, which can mislead readers’ initial interpretations. Semantics, on the other hand, while more closely related to high-level cognitive understanding, within the context of Linguistic Knowledge, pertains to the meanings of individual words or lexical semantics. This aspect is distinct from general conceptual knowledge, which falls under other dimensions of LLM capabilities, highlighting the meaning-related understanding, such as synonymy, antonymy, and hypernymy.
@@
Formal Knowledge. Beyond encoding linguistic structures and word meanings, an essential aspect of language capability involves understanding formal operations among words, or word-based reasoning. This means LLMs should be capable of recognizing relationships between words and deducing missing elements in a given pattern, such as completing analogies (e.g. “man:woman :: king:_”). FACE includes three types of reasoning mechanisms—deductive, inductive, and analogical reasoning—between words Bang et al. (2023). It also includes three categories of symbol-based formal skills: numeric (dealing with numbers), logic (applying logical operations), and manipulation (altering the inputs in a rule-based manner). An example task is concatenating the last letters of a word list (“think, machine, learning” “keg”).
@@
World Modeling. To step towards cognitive capabilities, well-grounded comprehension of factual and commonsense knowledge is required. Precisely, we decompose this capability into two primary mechanisms: remember and understand, respectively modeling the retrieval-based and comprehension-based capability Sugawara et al. (2020); Wang et al. (2022a). Considering the versatility of knowledge sources, we instantiate the remember sub-capability as recalling factual knowledge (open-ended facts), reading comprehension (facts in context), and applying commonsense reasoning. Based on the multiple granularities of text comprehension and hierarchy of input text, we characterize the understand sub-capability as two skills: understanding narrative or event structure (paragraph-level), and discourse comprehension (document-level).
@@
Social Modeling. The utility of human language lies in not only the understanding of the text itself but also the social context and mental states underlying communication, i.e. serving as a medium for information exchange between individuals. As suggested by research of cognitive science Adolphs (2009), the human brain has dedicated machinery, namely “theory of mind network”, for processing social information and understanding somebody’s mental state. Besides, since there are a lot of phenomena about non-literal language comprehension in daily life, such as jokes, sarcasm, and indirect speech, successful LLMs should be capable of applying social inference skills to attain the intended meaning beyond the literal content. In this paper, we incorporate two kinds of social modeling into FACE, encompassing pragmatics and theory of mind (ToM) reasoning. Pragmatics is evaluated by six kinds of dialogue, including polite deceits, irony, maxims of conversation, metaphor, indirect speech, and humor, while ToM is based on the “unexpected tasks” devised by Kosinski (2023).
2.2 FACE
@@
Based on the formulation of LLMs’ capabilities, we collect input and output pairs from various benchmarks and modify collected instances, yielding a unified question-answering (QA) format. After that, following the widely adopted few-shot in-context learning (ICL) Brown et al. (2020), we devise capability-specific instruction and frame FACE via instruction following Ouyang et al. (2022). We further leverage chain-of-thought (CoT) Wei et al. (2022a, b) style prompting to elicit two intermediate reasoning steps from the model, namely crystallized step and fluid step. The terms “crystallized” and “fluid” are borrowed from Cattel’s theory Cattell (1963), a foundational building block of cognitive science about the source of intelligence. Cattel’s theory delineates that crystallized intelligence is semantic knowledge from past experiences, fluid intelligence is the ability to navigate novel situations, and problem-solving uses both. Therefore, we add two intermediate steps to operationalize the two kinds of mechanisms, and aim to measure how well the model recalls and applies knowledge. Last, we compare the intermediate results with the reference answers to score the reasoning sub-steps, hence providing an assessment of the crystallized performance and fluid performance as well as problem-solving performance.
@@
As depicted in Figure 1, the pipeline FACE can be divided into three steps, including capability-specific instruction design, instruction decomposition, and stepwise evaluation. Specifically, we devise natural instruction for each capability-related task. Borrowing the widely used template schema of instruction-following Wang et al. (2022b); Mishra et al. (2022), the capability-specific instruction is comprised of three parts: title, task description and few-shot demonstrations. Precisely, the title defines a given QA task in high-level natural language and highlights the associated skills, while the task description not only presents a complete clarification of how an input text is expected to be mapped to final output, but also define the output of reasoning sub-steps through instruction decomposition. After that, following the given task description, a few in-context demonstrations are provided to better steer the response generation. At last, we collect the response results for each reasoning sub-steps, denoted as , and respectively evaluate them with the reference answer, denoted as , which are directly extracted or manually constructed from corresponding benchmarks. Formally, the procedure can be represented as:
@@
Instruction decomposition. Motivated by building explainable evaluation methods for machine reading comprehension (MRC) Ray Choudhury et al. (2022), which extracts relevant text spans and validates reasoning path, we propose to evaluate intermediate steps when LLMs apply a specific capability to solve practical problems. Specifically, we decompose the capability-specific instruction and add two intermediate steps, including crystallized and fluid steps. In practice, we leverage CoT-like iterative prompting strategy to elicit the intermediate reasoning from the model. Differing from the standard CoT that outputs a continuous rationale before the final answer, we first decompose the given question as follow-up sub-questions. After that, these sub-questions are used to help the model talk with itself to respectively discover (i) what knowledge this question is about, (ii) how to apply relevant knowledge to the given instance, and (iii) the final answer. In other words, FACE convert the CoT continuous rationale into easily parseable multi-step rationales, which externalizes reasoning of the model Shwartz et al. (2020); Zhou et al. (2022b) and enables the evaluation of crystallized performance and fluid performance. Formally, as depicted in the demonstrations of Figure 1, we expect that the the model outputs as: , where [Thought] can reason about the current situation, [Action] can be either (1) [Follow-up Question], which returns a sub-question, or (2) [Finish], and [Answer] is extracted as the reasoning result of a sub-step.
| Capability | Benchmark | QA type |
| Agreements | BLiMP Warstadt et al. (2020) | M |
| Licensing | Marvin and Linzen (2018) | M |
| Long-distance dependency | Wilcox et al. (2019) | M |
| Garden-path effects | Futrell et al. (2018) | M |
| Lexical semantics | Petersen and Potts (2023) | M |
| Deductive | Bang et al. (2023) | M |
| Inductive | Bang et al. (2023) | M |
| Analogical | Webb et al. (2023) | G |
| Numeric | MAWPS Koncel-Kedziorski et al. (2016) | G |
| Logic | Tian et al. (2021) | G |
| Manipulatation | Wei et al. (2022b) | G |
| Factual Knowledge | LAMA Petroni et al. (2019) | G |
| Reading Comprehension | Dua et al. (2019) | M |
| Commonsense | Talmor et al. (2019) | M |
| Discourse | Wang et al. (2023c) | G |
| Narrative | Xu et al. (2022) | G |
| Pragmatics | Hu et al. (2023) | M |
| Theory of mind | Ullman (2023) | G |

### v2

# FAC2E: Better Understanding Large Language Model Capabilities by Dissociating Language and Cognition

@@ Abstract
Large language models (LLMs) are primarily evaluated by overall performance on various text understanding and generation tasks. However, such a paradigm fails to comprehensively differentiate the fine-grained language and cognitive skills, rendering the lack of sufficient interpretation to LLMs’ capabilities. In this paper, we present FAC2E, a framework for Fine-grAined and Cognition-grounded LLMs’ Capability Evaluation. Specifically, we formulate LLMs’ evaluation in a multi-dimensional and explainable manner by dissociating the language-related capabilities and the cognition-related ones. Besides, through extracting the intermediate reasoning from LLMs, we further break down the process of applying a specific capability into three sub-steps: recalling relevant knowledge, utilizing knowledge, and solving problems. Finally, FAC2E evaluates each sub-step of each fine-grained capability, providing a two-faceted diagnosis for LLMs. Utilizing FAC2E, we identify a common shortfall in knowledge utilization among models and propose a straightforward, knowledge-enhanced method to mitigate this issue. Our results not only showcase promising performance enhancements but also highlight a direction for future LLM advancements.
FAC2E: Better Understanding Large Language Model Capabilities by Dissociating Language and Cognition
Xiaoqiang Wang1,2, Lingfei Wu3, Tengfei Ma4 and Bang Liu1,2 1DIRO & Institut Courtois, Université de Montréal 2Mila - Quebec AI Institute; 3Anytime.AI; 4Stony Brook University {xiaoqiang.wang, bang.liu}@umontreal.ca lwu@anytime-ai.com, tengfei.ma@stonybrook.edu
@@ 1 Introduction
To address this challenge, some studies have assessed the performance of LLMs on different tasks based on independent benchmarks from various dimensions Liang et al. (2022); Srivastava et al. (2023); Gao et al. (2023), such as commensense Zellers et al. (2019), knowledge Hendrycks et al. (2020); Yu et al. (2023), instruction-following Gu et al. (2024), and trustworthy Sun et al. (2024). Besides, motivated by building LLM-based AI assistants, other studies propose highly curated benchmarks with instance-level fine-grained annotations, such as difficulty and reasoning skills Mialon et al. (2023); Ye et al. (2024), for holistic evaluation of LLMs.
However, the existing studies, assessing effectiveness across various tasks, provided limited insight into the models’ true capabilities, as it only indicates their overall performance on specific datasets, without revealing the fine-grained capabilities acquired or their proficiency levels within the multiple capabilities involved. For instance, in the context of generative question answering, a model adept at extracting information but struggling to form a coherent understanding may exhibit similar overall performance to another model with profound understanding insights but difficulties in articulating accurate responses.
@@
We argue that a fine-grained understanding of LLMs’ capabilities can not only accurately unveil their inherent limitations, but also help us to better identify why one model outperforms the other and how the different capabilities correlate. Additionally, such insights allow us to provide tailored guidance to improve training efficiency or facilitate more advanced model development.
@@
In this paper, we propose FAC2E, a fine-grained capability evaluation framework for LLMs. Specifically, FAC2E dissociates the language-related and cognition-related capabilities of LLMs and organizing them into four distinct axes: Linguistic Knowledge, Formal Knowledge, World Modeling, and Social Modeling. This categorization is grounded in neuroscience evidence manifesting that language processing and cognitive processes, like memory and reasoning, operate differently in the brain. Drawing from this insight, we adapt a range of existing benchmarks into a unified question-answering format. We then develop specific instructions for each capability, allowing FAC2E to evaluate LLMs through a method known as few-shot instruction-following.
@@
Furthermore, we break down the application of a specific capability into three sub-steps: knowledge recall, knowledge utilization, and problem-solving, by iteratively drawing out the model’s intermediate reasoning. After evaluating each sub-step, FAC2E can reveal the quality of knowledge encoded in the model, and effectiveness in applying relevant knowledge to solve practical problems, offering a more comprehensive evaluation than a single performance metric could.
@@
Our findings reveal a notable gap in capabilities between open-source and proprietary models, especially for cognition-related capabilities. Additionally, we found that many models have difficulties in applying knowledge effectively. To address this, we suggest a knowledge-enhance remedy by incorporating relevant knowledge text as additional input. Experimental results show that it can help the backbone model (e.g. LLaMA 2) achieve approximately 90% of the performance of its instruction-tuned counterpart (e.g. LLaMA 2-Chat).
@@
| Conducting word-based formal reasoning through understanding lexical semantics. | Skill: | numeric, logic, and manipulation. | |
@@ 2 Methodology
In this section, we introduce FAC2E framework, designed for fine-grained and cognition-grounded LLMs’ capability evaluation. Specifically, we first define the taxonomy for LLMs’ capabilities based on the distinction between language and cognition, which is drawn upon insights from neuroscience Fedorenko and Varley (2016); Mahowald et al. (2023). Based on this, we transform a variety of existing benchmarks into the unified question-answering format, design capability-specific instruction, and frame FAC2E via few-shot instruction-following. Furthermore, we break down the evaluation process for each capability into a three-step reasoning approach. This involves identifying the knowledge pertinent to the input, examining how the model applies this knowledge in practical contexts, and assessing the effectiveness of its problem-solving. By evaluating each of these steps, FAC2E provides a comprehensive overview of the model’s performance, offering a more nuanced understanding of LLMs’ intrinsic capabilities.
@@ 2.1 Formulation of LLMs’ Capabilities
Human language processing, long studied in cognitive science and neuroscience, robustly attributes language and cognition to different brain areas, namely “language network” and “multi-demand network” Duncan (2010); Scott et al. (2017). The former is sensitive to linguistic regularities and formal operations, with damage leading to linguistic deficits, while the latter responds actively to various cognitively demanding processes, such as reasoning and memory. Similarly, prior analyses have identified core linguistic regions Zhang et al. (2024b) and language-independent knowledge neurons Chen et al. (2023) in LLMs, represented by different parameter sets and subnetworks, each contributing distinctly to language and reasoning tasks.
Motivated by this separated relationship, we define the LLMs’ capabilities as a 4-dimensional schema as shown in Table 1. Compared to the fine-grained definitions of high-level knowledge Hendrycks et al. (2020); Yu et al. (2023) and reasoning skills Ye et al. (2024) in related works, this formulation aims to dissociate language-related and cognition-related capabilities to define a broader range of both high-level and more fundamental low-level functionality. Additionally, it seeks to minimize the coupling between these capabilities to facilitate nuanced analysis (Sec. 3.1) and targeted improvement of the model (Sec. 3.2).
@@
Linguistic Knowledge. To effectively generate language, LLMs first understand text at a basic linguistic level, including both grammar and semantics. Grammaticality encompasses the rules that govern language structure, spanning from the sounds (phonological) and words (lexical) to the arrangement of words in phrases and sentences. To capture this grammatical structure as comprehensively as possible, especially given the challenges conventional models face in benchmarks like BLiMP Warstadt et al. (2020), we focus on four key skills: agreement (anaphora and subject-verb relationships), licensing (negative polarity items and reflexive pronouns), managing long-distance dependencies (filler-gap constructions and cleft sentences), and navigating garden-path sentences, which contain temporary ambiguities that must be resolved for correct understanding. For example, “the horse raced past the barn fell,” the initial interpretation is that the horse is racing, but upon reaching “fell,” it becomes clear that the horse is being raced by another (unnamed) entity.
Semantics, on the other hand, while more closely related to high-level cognitive understanding, within the context of Linguistic Knowledge, pertains to the meanings of individual words or lexical semantics Geeraerts (2009). This aspect is distinct from conceptual knowledge, which falls under other dimensions of LLM capabilities, highlighting the meaning-related understanding, such as synonymy, antonymy, and hypernymy.
@@
Formal Knowledge. Beyond encoding linguistic structures and word meanings, an essential aspect of language capability involves understanding formal operations among words, or word-based reasoning. This means LLMs should be capable of recognizing relationships between words and deducing missing elements in a given pattern, such as completing analogies (e.g. “man:woman :: king:_”). FAC2E includes three types of reasoning mechanisms—deductive, inductive, and analogical reasoning—between words Bang et al. (2023), and includes three symbol-based formal skills: numeric (dealing with numbers), logic (applying logical operations), and manipulation (altering the inputs in a rule-based manner) Wei et al. (2022b). An example task is concatenating the last letters of a word list (“think, machine, learning” “keg”).
@@
World Modeling. To step towards cognitive capabilities, well-grounded comprehension of factual and commonsense knowledge is required. Precisely, we decompose this capability into two primary mechanisms: remember and understand, respectively modeling the retrieval-based and comprehension-based capability Sugawara et al. (2020). Considering the versatility of knowledge sources, we instantiate the remember sub-capability as recalling factual knowledge (open-ended facts), reading comprehension (facts in context), and applying commonsense reasoning. Based on the multiple granularities of text comprehension and hierarchy of input text, we characterize the understand sub-capability as two skills: understanding narrative or event structure (paragraph-level), and discourse comprehension (document-level).
@@
Social Modeling. The utility of human language lies in not only the understanding of the text itself but also the social context and mental states underlying communication Adolphs (2009), i.e. serving as a medium for information exchange between individuals. Specifically, there are a lot of phenomena about non-literal language comprehension in daily life, such as jokes, sarcasm, and indirect speech, successful LLMs should be capable of applying social inference skills to attain the intended meaning beyond the literal content. In this paper, we incorporate two kinds of social modeling into FAC2E, encompassing pragmatics and theory of mind (ToM) reasoning. Pragmatics is evaluated by six kinds of dialogue, including polite deceits, irony, maxims of conversation, metaphor, indirect speech, and humor, while ToM is based on the “unexpected tasks” devised by Kosinski (2023).
2.2 FAC2E
@@
Based on the formulation of LLMs’ capabilities, we collect input and output pairs from various benchmarks and modify collected instances, yielding a unified question-answering (QA) format. After that, following the widely adopted few-shot in-context learning (ICL) Brown et al. (2020), we devise capability-specific instruction and frame FAC2E via instruction following Ouyang et al. (2022). We further leverage chain-of-thought (CoT) Wei et al. (2022a, b) style prompting to elicit two intermediate reasoning steps from the model, namely crystallized step and fluid step. The terms “crystallized” and “fluid” are borrowed from Cattel’s theory Cattell (1963), a foundational building block of cognitive science about the source of intelligence. Cattel’s theory delineates that crystallized intelligence is semantic knowledge from past experiences, fluid intelligence is the ability to navigate novel situations, and problem-solving uses both. Therefore, we add two intermediate steps to operationalize the two kinds of mechanisms, and aim to measure how well the model recalls and applies knowledge. Last, we compare the intermediate results with the reference answers to score the reasoning sub-steps, hence providing an assessment of the crystallized performance and fluid performance as well as problem-solving performance.
@@
As depicted in Figure 1, the pipeline FAC2E can be divided into three steps, including capability-specific instruction design, instruction decomposition, and stepwise evaluation. Specifically, we devise natural instruction for each capability-related task. Borrowing the widely used template schema of instruction-following Wang et al. (2022b); Mishra et al. (2022), the capability-specific instruction is comprised of three parts: title, task description and few-shot demonstrations. Precisely, the title defines a given QA task in high-level natural language and highlights the associated skills, while the task description not only presents a complete clarification of how an input text is expected to be mapped to final output, but also define the output of reasoning sub-steps through instruction decomposition. After that, following the given task description, a few in-context demonstrations are provided to better steer the response generation. At last, we collect the response results for each reasoning sub-steps, denoted as , and respectively evaluate them with the reference answer, denoted as , which are directly extracted from corresponding benchmarks. Formally, the procedure can be represented as:
@@
Instruction decomposition. We leverage CoT-like iterative prompting strategy to elicit the intermediate reasoning from the model to frame crystallized and fluid steps. Differing from the standard CoT that outputs a continuous rationale before the final answer, we first decompose the given question as follow-up sub-questions. After that, these sub-questions are used to help the model talk with itself to respectively discover (i) what knowledge this question is about, (ii) how to apply relevant knowledge to the given instance, and (iii) the final answer. In other words, FAC2E convert the CoT continuous rationale into easily parseable multi-step rationales, which externalizes reasoning of the model Shwartz et al. (2020); Zhou et al. (2022b) and enables the evaluation of crystallized performance and fluid performance. Formally, as depicted in Figure 1, we expect that the model outputs as: , where [Thought] can reason about the current situation, [Action] can be either (1) [Follow-up Question], which returns a sub-question, or (2) [Finish], and [Answer] is extracted as the reasoning result of a sub-step.
| Capability | Benchmark | #Samples | Length | QA |
| Agreements | BLiMP Warstadt et al. (2020) | 400 | 33 | M |
| Licensing | Marvin and Linzen (2018) | 1,500 | 36 | M |
| Long-distance dependency | Wilcox et al. (2019) | 660 | 48 | M |
| Garden-path effects | Futrell et al. (2018) | 450 | 40 | M |
| Lexical semantics | Petersen and Potts (2023) | 1,000 | 8 | M |
| Deductive | Bang et al. (2023) | 1,600 | 18 | M |
| Inductive | Bang et al. (2023) | 1,500 | 16 | M |
| Analogical | Webb et al. (2023) | 800 | 4 | G |
| Numeric | MAWPS Koncel-Kedziorski et al. (2016) | 400 | 33 | G |
| Logic | Tian et al. (2021) | 1,000 | 87 | G |
| Manipulatation | Wei et al. (2022b) | 250 | 31 | G |
| Factual Knowledge | LAMA Petroni et al. (2019) | 500 | 48 | G |
| Reading Comprehension | Dua et al. (2019) | 1,000 | 195 | M |
| Commonsense | Talmor et al. (2019) | 800 | 13 | M |
| Discourse | Wang et al. (2023c) | 400 | 439 | G |
| Narrative | Xu et al. (2022) | 200 | 395 | G |
| Pragmatics | Hu et al. (2023) | 150 | 288 | M |
| Theory of mind | Ullman (2023) | 80 | 152 | G |